OpenStudy (anonymous):

Identify the vertex for the graph of y = -3x2 - 6x + 4.

OpenStudy (mertsj):

The x coordinate of the vertex is -b/2a. In this case that is 6/-6 or -1. Now find the y coordinate of the vertex by plugging -1 into the given equation and finding y: -3(-1)^2-6(-1)+4= -3+6+4=7 So the vertex is (-1,7)